A block slides from rest from the top of an inclined plane 10 m long which is inclined 37° with



the horizontal. If the coefficient of kinetic friction is 0.30, determine how long it will take the



block to reach the bottom of the plane.

1
Expert's answer
2022-01-27T10:14:07-0500

Given:

"l=10\rm \: m"

"\theta=37^\circ"

"\mu_k=0.30"


The Newton's second law says

"a=\frac{F_{\rm net}}{m}=\frac{mg\sin\theta-\mu_k mg\cos\theta}{m}"

So, the acceleration of the block

"a=g(\sin\theta-\mu_k \cos\theta)"

Hence, time of motion

"t=\sqrt{2l/a}=\sqrt{2l/g(\sin\theta-\mu_k \cos\theta)}"

"t=\sqrt{2*10/9.8(\sin37^\circ-0.30\cos 37^\circ)}=2.37\:\rm s"